a fxn is defined on[-1,1]and
area of sq. is 4.
with its vertices at (0,0) and (x,f(x))
then f(x) is...


Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Area of Square: 4
=>The sides of square are of 2 units each.
Diagonal Length:
\sqrt{(f(x))^{2}+x^{2}} = 2\sqrt{2}
(f(x))^{2}+x^{2}} = 8
f(x) = \pm \sqrt{8-x^{2}}
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty